# Compatible Relative Lefschetz Fibrations on Admissible Relative Stein Pairs

### v2

@@ 3. Suspended Legendrian 2-links in the complement of Brieskorn manifolds
For the purpose of the present paper, one needs to define suspendibility of a relative Legendrian $2$-link $(L,K)$ embedded in the contact boundary of a subcritical Stein $6$-domain. If $\widetilde{W}$ is such a domain, then it can be written as a product $\widetilde{W}=\widetilde{X}\times D^{2}$ for some Stein $4$-domain $\widetilde{X}$, and the pair $(\widetilde{W},\widetilde{X})$ is a relative Stein pair because it clearly satisfies the conditions (0)-(2) of Definition 1.1. (The Stein handles of $\widetilde{W}$ can be taken as the thickening of those of $\widetilde{X}$). Note that using a given Stein handle decomposition of $\widetilde{X}$, $(\widetilde{W},\widetilde{X})$ can be described by a relative Stein diagram in the dotted form containing only dotted unknots ($1$-handles of $\widetilde{X}$, and so $\widetilde{W}$) and the framed dashed knots ($2$-handles of $\widetilde{X}$, and so $\widetilde{W}$). One should note that such a diagram is not unique since one can start with a different Stein handle decomposition of $\widetilde{X}$ (still defining the same Stein structure on $\widetilde{X}$ (and so on $\widetilde{W}$) up to Stein deformation). So, let $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$ be the relative Stein diagram (in the dotted form) of $\widetilde{W}=\widetilde{X}\times D^{2}$ obtained from a given Stein handle decomposition $\mathcal{H}$ of $\widetilde{X}$. Recall that, in a diagram $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$, replacing dotted unknots with $(+1)$-framed unknots, we obtain a relative contact surgery diagram, denoted by $CSD_{\mathcal{H}}(\partial(\widetilde{X}\times D^{2}))$, describing the contact boundary $\partial(\widetilde{X}\times D^{2})$. Therefore, one can realize $(L,K)$ in this surgery diagram by drawing the components of $K$ as solid knots. Forgetting framings on the $(+1)$-framed unknots and the framed dashed knots in the diagram, we obtain a Legendrian $1$-link $K_{\mathcal{H}}$ (in the standard contact $S^{3}$) which consists of the above (unframed) unknots and (unframed) dashed knots. Note that after forgetting all the framings in the diagram $CSD_{\mathcal{H}}(\partial(\widetilde{X}\times D^{2}))$, the link $(L,K)$ becomes a relative Legendrian $2$-link in the standard contact pair $(S^{5},S^{3})$.
@@
A relative Legendrian $2$-link $(L,K)$ in the contact boundary $\partial(\widetilde{X}\times D^{2})$ of a subcritical Stein $6$-domain $\widetilde{W}=\widetilde{X}\times D^{2}$ is said to be suspendible if there exists a Stein handle decomposition $\mathcal{H}$ of $\widetilde{X}$ such that there exists a Legendrian isotopy in the complement $S^{5}\setminus K_{\mathcal{H}}$ from $(L,K)$ to a suspended Legendrian $2$-link based on some Legendrian $1$-link $K^{\prime}$ (possibly different than $K$) in the complement $S^{3}\setminus K_{\mathcal{H}}$.
@@
which splits as $W_{2}=\widetilde{X}\times D^{2}$ where $\widetilde{X}$ is a Stein $4$-domain. By assumption $(L,K)\subset\partial(\widetilde{X}\times D^{2})$ is suspendible, therefore, there exists a Stein handle decomposition $\mathcal{H}$ of $\widetilde{X}$ such that we can isotope $(L,K)$ (through Legendrian $2$-links in $S^{5}\setminus K_{\mathcal{H}}$) to some suspended Legendrian $2$-link. Assuming Legendrian isotopy taking $(L,K)$ to a suspended Legendrian $2$-link has been already performed, we may assume that $(L,K)$ is suspended.
@@
Every admissible Stein pair without relative Stein $2$-handles admits a compatible relative Lefschetz fibration.
@@
The relative Stein diagram of an admissible Stein pair $(W,X)$ without relative $2$-handles consists of $r$ dotted unknots $U_{1},...,U_{r}$ (representing relative $1$-handles) and the solid knots drawn for each component of $K$ (the base of the suspended Legendrian $2$-link $(L,K)$ along which the relative $3$-handles are attached) as discussed above. (Here, in particular, it is assumed that this relative Stein diagram of $(W,X)$ is obtained from a diagram $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$ for a suitable Stein handle decomposition $\mathcal{H}$.) First, we forget the dots on $U_{i}$’s and consider the entire diagram as a relative $2$-link in the relative standard contact pair $(S^{5},S^{3})$. Then, by Proposition 3.1, there exists a pair of Brieskorn manifolds $(\Sigma(p,q,2),\Sigma(p,q))$ in $(S^{5},S^{3})$ such that $(L,K)$ is properly embedded on their Seifert fillings $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ as Legendrian. Note that each $U_{i}$ is also realized on $V_{\epsilon}(p,q)\subset V_{\epsilon}(p,q,2)$. Moreover, by Proposition 2.1, there exist Lefschetz fibrations on $D^{6}$ and $D^{4}$ with regular fibers $F_{D^{6}}=V_{\epsilon}(p,q,2)$ and $F_{D^{4}}=V_{\epsilon}(p,q)$, respectively, as depicted in Figure 6-(a). Recall (from Remark 2.3) that the monodromies of these fibrations are denoted by $\widetilde{\varphi}_{p,q}$ and $\varphi_{p,q}$, respectively, which are the product of $(p-1)(q-1)$ many right-handed Dehn twists along explicit spheres.
@@
First, isotope each $U_{i}$ near the binding $\Sigma(p,q)$ of the induced open book on the boundary $S^{3}=\partial D^{4}$ as in [1] so that it becomes transverse to each page $F_{D^{4}}$. Once this is achieved, $U_{i}$ will also transversally intersect (near the binding $\Sigma(p,q,2)$) each page $F_{D^{6}}$ of the induced open book on the boundary $S^{5}=\partial D^{6}$ (thanks to the relative Lefschetz fibration structure). For each $i=1,...,r$, attaching a $1$-handle to $D^{4}$ corresponds to first pushing the interior of the spanning disk of $U_{i}$ into $D^{4}$ and then removing a tubular neighborhood of this resulting disk. This is equivalent to removing a small $2$-disk from each fiber $F_{D^{4}}$. Since each fiber $F_{D^{4}}$ is properly embedded in the corresponding fiber $F_{D^{6}}$, the tubular neighborhood of this small $2$-disk in the fiber $F_{D^{6}}$ is a small $4$-disk. On the other hand, by the condition (1) in Definition 1.1, the cocore of each $1$-handle of $X$ is properly embedded in the cocore of the corresponding $1$-handle of $W$. Therefore, simultaneously removing these small $4$-disks from the fibers $F_{D^{6}}$ corresponds to attaching the corresponding $6$-dimensional relative $1$-handle of $W$ to $D^{6}$. In other words, attaching each relative Stein $1$-handle of $(W,X)$ to the Stein pair $(D^{6},D^{4})$ is equivalent to removing a small disk pair $(D^{4},D^{2})$ from each fiber pair $(F_{D^{6}},F_{D^{4}})=(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$. This is depicted in Figure 6-(b) where a single relative $1$-handle is considered (for simplicity), and the “punctured” fibers (i.e., new fibers after removing disks) are denoted by $\mathring{F}_{D^{6}}$ and $\mathring{F}_{D^{4}}$.
@@
After attaching all relative Stein $1$-handles of $(W,X)$ to $(D^{6},D^{4})$, we get a new compatible relative Lefschetz fibration on $(D^{6}\cup 1$-handles, $D^{4}\cup 1$-handles)$=(\natural_{r}S^{1}\times D^{5},\natural_{r}S^{1}\times D^{3})$ whose regular fibers are $r$ times “punctured” $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ near boundaries, which we denote by
$(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})=(\mathring{V}_{\epsilon}(p,q,2),\mathring{V}_{\epsilon}(p,q)),$
and whose monodromy is obtained by extending $\widetilde{\varphi}_{p,q}$ and $\varphi_{p,q}$ (over the $1$-handles) by identity.
Now assume that $(L,K)$ has $\ell$ components, say $(L_{j},K_{j})$, $j=1,...,\ell$. So, for each $j$, we attach a relative Stein $3$-handle $({H}_{j},{h}_{j})$ to $(\natural_{r}S^{1}\times D^{5},\natural_{r}S^{1}\times D^{3})$ along the relative Legendrian $2$-knot $(L_{j},K_{j})$ which is properly embedded in $(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})$. Each ${H}_{j}$ (resp. ${h}_{j}$) is attached along a Legendrian $L_{j}$ (resp. $K_{j}$) which lies on a Weinstein (so symplectic) page $\mathring{F}_{D^{6}}$ (resp. $\mathring{F}_{D^{4}}$) of the boundary (compatible) open book so that the contact framing of $L_{j}$ (resp. $K_{j}$) coincides with its page framing on $\mathring{F}_{D^{6}}$ (resp. on $\mathring{F}_{D^{4}}$). Furthermore, recall that any Legendrian submanifold sitting on a page of a compatible open book is Lagrangian. Thus, we conclude that ${H}_{j}$ (resp. ${h}_{j}$) is a six (resp. four) dimensional Lefschetz handle. Consequently, attaching each relative Stein $3$-handle corresponds to attaching a “relative” Lefschetz handle. So, attaching all such handles constructs a compatible relative Lefschetz fibration, say $(\pi_{W},\pi_{X})$, on the admissible Stein pair $(W,X)$. Note that the fibers of $(\pi_{W},\pi_{X})$ are still $(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})$. On the other hand, the monodromy of $(\pi_{W},\pi_{X})$ is the pair
@@
On every subcritical Stein $6$-domain (no handles of critical index) there exists a relative Stein pair structure which admits a nontrivial compatible relative Lefschetz fibration.
@@
Let $\widetilde{W}$ be a subcritical Stein $6$-domain. By the result of Cieliebak [4], there is a splitting $\widetilde{W}=\widetilde{X}\times D^{2}$ where $\widetilde{X}$ is a Stein $4$-domain. Starting from a Stein handle decomposition $\mathcal{H}$ of $\widetilde{X}$, one obtains a compatible Lefschetz fibration on $\widetilde{X}$ with fibers $r$ times “punctured” $V_{\epsilon}(p,q)$ for some $p,q$ (where $r$ is the number of $1$-handles of $\widetilde{X}$ in $\mathcal{H}$), and whose monodromy is that of the torus link $\Sigma(p,q)$ composed with right-handed Dehn twists corresponding to the Stein (Lefschetz) $2$-handles of $\widetilde{X}$ in $\mathcal{H}$, as constructed in [1]. So, we have a (Lefschetz) fibration map $\pi_{\mathcal{H}}:\widetilde{X}\rightarrow D^{2}$ which has finitely many nondegenerate critical points.
@@
${\partial\widetilde{\pi}_{\mathcal{H}}\over\partial z_{0}}={\partial{\pi}_{\mathcal{H}}\over\partial z_{0}}=0,\quad{\partial\widetilde{\pi}_{\mathcal{H}}\over\partial z_{1}}={\partial{\pi}_{\mathcal{H}}\over\partial z_{1}}=0,\quad{\partial\widetilde{\pi}_{\mathcal{H}}\over\partial z_{2}}=0$
where $(z_{0},z_{1})$ are the local complex coordinates around a critical point in $\widetilde{X}$. So, the critical points of $\widetilde{\pi}_{\mathcal{H}}$ are the same as those of $\pi_{\mathcal{H}}$ with the additional zero $z_{2}$-coordinate. Also, if we check the Hessian of $\widetilde{\pi}_{\mathcal{H}}$ at each critical point, we see that it is nondegenerate because the corresponding Hessian of $\pi_{\mathcal{H}}$ is nondegenerate. Each regular fiber of $\widetilde{\pi}_{\mathcal{H}}$ is the $r$ times “punctured” Brieskorn variety $V_{\epsilon}(p,q,2)$, “suspension” of the corresponding fiber of $\pi_{\mathcal{H}}$, and the monodromy of $\widetilde{\pi}_{\mathcal{H}}$ is the product of right-handed Dehn twists along the spheres obtained by taking the suspensions of the corresponding vanishing cycles of $\pi_{\mathcal{H}}$. Therefore, $\widetilde{\pi}_{\mathcal{H}}$ is a compatible Lefschetz fibration map on $\widetilde{W}$ with nontrivial monodromy.
@@
where $\widetilde{W}$ is the subcritical part of $W$ which split as $\widetilde{W}=\widetilde{X}\times D^{2}$ for some Stein $4$-domain $\widetilde{X}$, and $(H_{j},h_{j})$ are the relative Stein $3$-handles of $W$ attached along a relative Legendrian $2$-link $(L,K)$. By Proposition 4.3, there exists a compatible relative Lefschetz fibration $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$ on the subcritical part $(\widetilde{W},\widetilde{X})$ of $(W,X)$ for any Stein handle decomposition $\mathcal{H}$ of $\widetilde{X}$. We choose $\mathcal{H}$ in such a way that the link $(L,K)$ is Legendrian isotopic (in $S^{5}\setminus K_{\mathcal{H}}$) to a suspended Legendrian $2$-link. Such a decomposition $\mathcal{H}$ exists by admissibility (i.e., by suspendibility of $(L,K)$, see Definition 3.6). Hence, assuming the above Legendrian isotopy has been already performed, we may assume $(L,K)$ is a suspended Legendrian $2$-link sitting (as Lagrangian) on a relative page $(\widetilde{F}_{\mathcal{H}},F_{\mathcal{H}})$ of the relative compatible open book (induced by $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$) on the boundary relative contact pair $(\partial\widetilde{W},\partial\widetilde{X})$. Note that by Lemma 3.2 the contact and the page framing of each component $(L_{j},K_{j})$ of $(L,K)$ coincide.
To finish the proof, one just needs to proceed as in the last part of the proof of Proposition 4.2. More precisely, we have observed that each relative Stein $3$-handle $(H_{j},h_{j})$ attached along $(L_{j},K_{j})$ is a relative Lefschetz handle, i.e., $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$ extends over each $(H_{j},h_{j})$. Thus, attaching all $(H_{j},h_{j})$’s, we get a compatible relative Lefschetz fibration, say $(\pi_{W},\pi_{X})$, on the admissible Stein pair $(W,X)$ with relative regular fibers $(\widetilde{F}_{\mathcal{H}},F_{\mathcal{H}})$. If $(\widetilde{\varphi}_{\mathcal{H}},\varphi_{\mathcal{H}})$ denotes the monodromy of $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$, then the monodromy of $(\pi_{W},\pi_{X})$ is the pair

### v3

@@ 3. Suspended Legendrian 2-links in the complement of Brieskorn manifolds
For the purpose of the present paper, one needs to define suspendibility of a relative Legendrian $2$-link $(L,K)$ embedded in the contact boundary of a subcritical Stein $6$-domain. If $\widetilde{W}$ is such a domain, then it can be written as a product $\widetilde{W}=\widetilde{X}\times D^{2}$ for some Stein $4$-domain $\widetilde{X}$, and the pair $(\widetilde{W},\widetilde{X})$ is a relative Stein pair because it clearly satisfies the conditions (0)-(2) of Definition 1.1. (The Weinstein handles of $\widetilde{W}$ can be taken as the thickening of those of $\widetilde{X}$). Note that using a given Weinstein handle decomposition of $\widetilde{X}$, $(\widetilde{W},\widetilde{X})$ can be described by a relative Stein diagram in the dotted form containing only dotted unknots ($1$-handles of $\widetilde{X}$, and so $\widetilde{W}$) and the framed dashed knots ($2$-handles of $\widetilde{X}$, and so $\widetilde{W}$). One should note that such a diagram is not unique since one can start with a different Weinstein handle decomposition of $\widetilde{X}$ (still defining the same Stein structure on $\widetilde{X}$ (and so on $\widetilde{W}$) up to Stein deformation). So, let $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$ be the relative Stein diagram (in the dotted form) of $\widetilde{W}=\widetilde{X}\times D^{2}$ obtained from a given Weinstein handle decomposition $\mathcal{H}$ of $\widetilde{X}$. Recall that, in a diagram $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$, replacing dotted unknots with $(+1)$-framed unknots, we obtain a relative contact surgery diagram, denoted by $CSD_{\mathcal{H}}(\partial(\widetilde{X}\times D^{2}))$, describing the contact boundary $\partial(\widetilde{X}\times D^{2})$. Therefore, one can realize $(L,K)$ in this surgery diagram by drawing the components of $K$ as solid knots. Forgetting framings on the $(+1)$-framed unknots and the framed dashed knots in the diagram, we obtain a Legendrian $1$-link $K_{\mathcal{H}}$ (in the standard contact $S^{3}$) which consists of the above (unframed) unknots and (unframed) dashed knots. Note that after forgetting all the framings in the diagram $CSD_{\mathcal{H}}(\partial(\widetilde{X}\times D^{2}))$, the link $(L,K)$ becomes a relative Legendrian $2$-link in the standard contact pair $(S^{5},S^{3})$.
@@
A relative Legendrian $2$-link $(L,K)$ in the contact boundary $\partial(\widetilde{X}\times D^{2})$ of a subcritical Stein $6$-domain $\widetilde{W}=\widetilde{X}\times D^{2}$ is said to be suspendible if there exists a Weinstein handle decomposition $\mathcal{H}$ of $\widetilde{X}$ such that there exists a Legendrian isotopy in the complement $S^{5}\setminus K_{\mathcal{H}}$ from $(L,K)$ to a suspended Legendrian $2$-link based on some Legendrian $1$-link $K^{\prime}$ (possibly different than $K$) in the complement $S^{3}\setminus K_{\mathcal{H}}$.
@@
which splits as $W_{2}=\widetilde{X}\times D^{2}$ where $\widetilde{X}$ is a Stein $4$-domain. By assumption $(L,K)\subset\partial(\widetilde{X}\times D^{2})$ is suspendible, therefore, there exists a Weinstein handle decomposition $\mathcal{H}$ of $\widetilde{X}$ such that we can isotope $(L,K)$ (through Legendrian $2$-links in $S^{5}\setminus K_{\mathcal{H}}$) to some suspended Legendrian $2$-link. Assuming Legendrian isotopy taking $(L,K)$ to a suspended Legendrian $2$-link has been already performed, we may assume that $(L,K)$ is suspended.
@@
Every admissible Stein pair without relative $2$-handles admits a compatible relative Lefschetz fibration.
@@
The relative Stein diagram of an admissible Stein pair $(W,X)$ without relative $2$-handles consists of $r$ dotted unknots $U_{1},...,U_{r}$ (representing relative $1$-handles) and the solid knots drawn for each component of $K$ (the base of the suspended Legendrian $2$-link $(L,K)$ along which the relative $3$-handles are attached) as discussed above. (Here, in particular, it is assumed that this relative Stein diagram of $(W,X)$ is obtained from a diagram $RSD_{\mathcal{H}}(\widetilde{X}\times D^{2})$ for a suitable Weinstein handle decomposition $\mathcal{H}$.) First, we forget the dots on $U_{i}$’s and consider the entire diagram as a relative $2$-link in the relative standard contact pair $(S^{5},S^{3})$. Then, by Proposition 3.1, there exists a pair of Brieskorn manifolds $(\Sigma(p,q,2),\Sigma(p,q))$ in $(S^{5},S^{3})$ such that $(L,K)$ is properly embedded on their Seifert fillings $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ as Legendrian. Note that each $U_{i}$ is also realized on $V_{\epsilon}(p,q)\subset V_{\epsilon}(p,q,2)$. Moreover, by Proposition 2.1, there exist Lefschetz fibrations on $D^{6}$ and $D^{4}$ with regular fibers $F_{D^{6}}=V_{\epsilon}(p,q,2)$ and $F_{D^{4}}=V_{\epsilon}(p,q)$, respectively, as depicted in Figure 5-(a). Recall (from Remark 2.3) that the monodromies of these fibrations are denoted by $\widetilde{\varphi}_{p,q}$ and $\varphi_{p,q}$, respectively, which are the product of $(p-1)(q-1)$ many right-handed Dehn twists along explicit spheres.
@@
First, isotope each $U_{i}$ near the binding $\Sigma(p,q)$ of the induced open book on the boundary $S^{3}=\partial D^{4}$ as in [1] so that it becomes transverse to each page $F_{D^{4}}$. Once this is achieved, $U_{i}$ will also transversally intersect (near the binding $\Sigma(p,q,2)$) each page $F_{D^{6}}$ of the induced open book on the boundary $S^{5}=\partial D^{6}$ (thanks to the relative Lefschetz fibration structure). For each $i=1,...,r$, attaching a $1$-handle to $D^{4}$ corresponds to first pushing the interior of the Lagrangian spanning disk of $U_{i}$ into $D^{4}$ and then removing a tubular neighborhood of this resulting disk. This is equivalent to puncture each fiber $F_{D^{4}}$. Since each fiber $F_{D^{4}}$ is properly embedded in the corresponding fiber $F_{D^{6}}$, this will correspond to carving a small neatly embedded Lagrangian $2$-disk in the fiber $F_{D^{6}}$. On the other hand, by the condition (1) in Definition 1.1, the cocore of each $1$-handle of $X$ is properly embedded in the cocore of the corresponding $1$-handle of $W$. Therefore, simultaneously carving these small neatly embedded Lagrangian $2$-disks from the fibers $F_{D^{6}}$ corresponds to attaching the corresponding $6$-dimensional relative $1$-handle of $W$ to $D^{6}$. In other words, attaching each relative $1$-handle of $(W,X)$ to the Stein pair $(D^{6},D^{4})$ is equivalent to carving a small Lagrangian disk pair $(D^{2},D^{0})$ from each fiber pair $(F_{D^{6}},F_{D^{4}})=(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$. This is depicted in Figure 5-(b) where a single relative $1$-handle is considered (for simplicity), and the “punctured” fibers (i.e., new fibers after carving disks) are denoted by $\mathring{F}_{D^{6}}$ and $\mathring{F}_{D^{4}}$.
After attaching all relative $1$-handles of $(W,X)$ to $(D^{6},D^{4})$, we get a new compatible relative Lefschetz fibration on $(D^{6}\cup 1$-handles, $D^{4}\cup 1$-handles)$=(\natural_{r}S^{1}\times D^{5},\natural_{r}S^{1}\times D^{3})$ whose regular fibers are $r$ times “punctured” $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ near boundaries, which we denote by $(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})=(\mathring{V}_{\epsilon}(p,q,2),\mathring{V}_{\epsilon}(p,q)),$ and whose monodromy is obtained by extending $\widetilde{\varphi}_{p,q}$ and $\varphi_{p,q}$ (over the $1$-handles) by identity.
@@
Now assume that $(L,K)$ has $\ell$ components, say $(L_{j},K_{j})$, $j=1,...,\ell$. So, for each $j$, we attach a relative $3$-handle $({H}_{j},{h}_{j})$ to $(\natural_{r}S^{1}\times D^{5},\natural_{r}S^{1}\times D^{3})$ along the relative Legendrian $2$-knot $(L_{j},K_{j})$ which is properly embedded in $(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})$. Each ${H}_{j}$ (resp. ${h}_{j}$) is attached along a Legendrian $L_{j}$ (resp. $K_{j}$) which lies on a Weinstein (so symplectic) page $\mathring{F}_{D^{6}}$ (resp. $\mathring{F}_{D^{4}}$) of the boundary (compatible) open book so that the contact framing of $L_{j}$ (resp. $K_{j}$) coincides with its page framing on $\mathring{F}_{D^{6}}$ (resp. on $\mathring{F}_{D^{4}}$). Furthermore, recall that any Legendrian submanifold sitting on a page of a compatible open book is Lagrangian. Thus, we conclude that ${H}_{j}$ (resp. ${h}_{j}$) is a six (resp. four) dimensional Lefschetz handle. Consequently, attaching each relative $3$-handle corresponds to attaching a “relative” Lefschetz handle. So, attaching all such handles constructs a compatible relative Lefschetz fibration, say $(\pi_{W},\pi_{X})$, on the admissible Stein pair $(W,X)$. Note that the fibers of $(\pi_{W},\pi_{X})$ are still $(\mathring{F}_{D^{6}},\mathring{F}_{D^{4}})$. On the other hand, the monodromy of $(\pi_{W},\pi_{X})$ is the pair
@@
On every subcritical Stein $6$-domain (no handles of critical index) there exists a relative Stein pair structure which admits a nontrivial compatible relative Lefschetz fibration with regular fibers $\mathring{V}_{\epsilon}(p,q,2)\cup\bigcup\limits_{j=1}^{2k}\tilde{h_{j}}$ for some $p,q$, where each $\tilde{h_{j}}$ is a $4$-dimensional $2$-handle.
@@
Let $\widetilde{W}$ be a subcritical Stein $6$-domain. By the result of Cieliebak [4], there is a splitting $\widetilde{W}=\widetilde{X}\times D^{2}$ where $\widetilde{X}$ is a Stein $4$-domain. Starting from a Weinstein handle decomposition $\mathcal{H}$ of $\widetilde{X}$, one obtains a compatible Lefschetz fibration on $\widetilde{X}$ with fibers $r$ times “punctured” $V_{\epsilon}(p,q)$ for some $p,q$ (where $r$ is the number of $1$-handles of $\widetilde{X}$ in $\mathcal{H}$), and whose monodromy is that of the torus link $\Sigma(p,q)$ composed with right-handed Dehn twists corresponding to the Stein (Lefschetz) $2$-handles of $\widetilde{X}$ in $\mathcal{H}$, as constructed in [1]. So, we have a (Lefschetz) fibration map $\pi_{\mathcal{H}}:\widetilde{X}\rightarrow D^{2}$ which has finitely many nondegenerate critical points.
@@
${\frac{\partial\widetilde{\pi}_{\mathcal{H}}}{\partial z_{0}}}={\frac{\partial{\pi}_{\mathcal{H}}}{\partial z_{0}}}=0,\quad{\frac{\partial\widetilde{\pi}_{\mathcal{H}}}{\partial z_{1}}}={\frac{\partial{\pi}_{\mathcal{H}}}{\partial z_{1}}}=0,\quad{\frac{\partial\widetilde{\pi}_{\mathcal{H}}}{\partial z_{2}}}=0$
where $(z_{0},z_{1})$ are the local complex coordinates around a critical point in $\widetilde{X}$. So, the critical points of $\widetilde{\pi}_{\mathcal{H}}$ are the same as those of $\pi_{\mathcal{H}}$ with the additional zero $z_{2}$-coordinate. Also, if we check the Hessian of $\widetilde{\pi}_{\mathcal{H}}$ at each critical point, we see that it is nondegenerate because the corresponding Hessian of $\pi_{\mathcal{H}}$ is nondegenerate. Each regular fiber of $\widetilde{\pi}_{\mathcal{H}}$ can be seen as double branched cover of $\widetilde{X}$ which is branched over ${\pi}_{\mathcal{H}}^{-1}(c)$ where $c$ is a regular value of $\widetilde{\pi}_{\mathcal{H}}$. With this point of view, the effect of attaching $k$-many $4$-dimensional $2$-handles to the fibers of $\widetilde{\pi}_{\mathcal{H}}$ will result in attaching $2k$-many $2$-handles to $r$-times punctured $V_{\epsilon}(p,q,2)$. The monodromy of $\widetilde{\pi}_{\mathcal{H}}$ is the product of right-handed Dehn twists along the spheres obtained by taking the suspensions of the corresponding vanishing cycles of $\pi_{\mathcal{H}}$. Therefore, $\widetilde{\pi}_{\mathcal{H}}$ is a compatible Lefschetz fibration map on $\widetilde{W}$ with nontrivial monodromy.
@@
where $\widetilde{W}$ is the subcritical part of $W$ which split as $\widetilde{W}=\widetilde{X}\times D^{2}$ for some Stein $4$-domain $\widetilde{X}$, and $(H_{j},h_{j})$ are the relative $3$-handles of $W$ attached along a relative Legendrian $2$-link $(L,K)$. By Proposition 4.3, there exists a compatible relative Lefschetz fibration $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$ on the subcritical part $(\widetilde{W},\widetilde{X})$ of $(W,X)$ for any Weinstein handle decomposition $\mathcal{H}$ of $\widetilde{X}$. We choose $\mathcal{H}$ in such a way that the link $(L,K)$ is Legendrian isotopic (in $S^{5}\setminus K_{\mathcal{H}}$) to a suspended Legendrian $2$-link. Such a decomposition $\mathcal{H}$ exists by admissibility (i.e., by suspendibility of $(L,K)$, see Definition 3.6). Hence, assuming the above Legendrian isotopy has been already performed, we may assume $(L,K)$ is a suspended Legendrian $2$-link sitting (as Lagrangian) on a relative page $(\widetilde{F}_{\mathcal{H}},F_{\mathcal{H}})$ of the relative compatible open book (induced by $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$) on the boundary relative contact pair $(\partial\widetilde{W},\partial\widetilde{X})$. Note that by Lemma 3.2 the contact and the page framing of each component $(L_{j},K_{j})$ of $(L,K)$ coincide.
To finish the proof, one just needs to proceed as in the last part of the proof of Proposition 4.2. More precisely, we have observed that each relative $3$-handle $(H_{j},h_{j})$ attached along $(L_{j},K_{j})$ is a relative Lefschetz handle, i.e., $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$ extends over each $(H_{j},h_{j})$. Thus, attaching all $(H_{j},h_{j})$’s, we get a compatible relative Lefschetz fibration, say $(\pi_{W},\pi_{X})$, on the admissible Stein pair $(W,X)$ with relative regular fibers $(\widetilde{F}_{\mathcal{H}},F_{\mathcal{H}})$. If $(\widetilde{\varphi}_{\mathcal{H}},\varphi_{\mathcal{H}})$ denotes the monodromy of $(\widetilde{\pi}_{\mathcal{H}},\pi_{\mathcal{H}})$, then the monodromy of $(\pi_{W},\pi_{X})$ is the pair